Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?
Answer by Paul(988)

Let the number of dimes be x
Let the number of nickels be y
Nickel value times 100 = 5
Dime value times 100 = 10
5.65 times 100 = 565
10.45 times 100 = 1045

5x+10y=565
5x=565-10y
x=113-2y (subsituition)

Dimes doubled = 2x
Nickles increased by 8 = y+8
10(2x)+5(y+8)=1045
20x+5y+40=1045
Subsitute for x
20(113-2y)+5y=1005
2260-40y+5y=1005
-35y=-1255
y=36 (rounded)
x=41

Hence, there are 36 dimes and 41 nickels.
Paul.